RawlsK29:

How many solutions can be found for the equation 4z + 2(z − 4) = 3z + 11? None <-- I think this one One <-- or maybe this one Two Infinitely many

RawlsK29:

@Warriorz13 @Shadow

Shadow:

\[4z + 2 (z - 4) = 3z + 11\] Distributive Property \[4z + 2z - 8 = 3z + 11\] Like terms \[6z - 8 = 3z + 11\] Add 8 to both sides Subtract 3z from both sides \[3z = 19\] \[x = \frac{ 19 }{ 3 }\]

Shadow:

What does that look like to you? :)

RawlsK29:

Umm... one solution?

Shadow:

Yes. Two solutions would be if we had plus or minus some number Infinite solutions would look like 3=3 or x = x No solutions would look like 3 = 5

RawlsK29:

Oh okay, thank you!😎